Burger Inc. has a market capitalization of $100 million, and debt outstanding of $40 million. Burger Inc. plans to maintain tplease show all workYear End Dividends Paid* S&P 500 Realized Return Microsoft Realized Return 1-Month T-Bill Return 2004 2005 2006 2007 2008 200
Answer #1

1. WACC

We know, value of firm can be calculated as below

Value of firm= Free Cash Flow expected/ Wacc- Growth

100= 7/ (WACC-3%)

WACC-3% = 7/100

WACC= 7% + 3%

WACC= 10%

2. Interest Tax Shield= Interest Paid * Rate of Tax

= 40*7.5%*21*

=0.63M

Interest tax shield is the tax shield available with the borrower for the interest paid on debt. It eventually reduces the cost of debt for the borrower.